# FilterPrompt: Guiding Image Transfer in Diffusion Models

### v2

@@ 4.1 FilterPrompt
Given the dynamic and evolving nature of the diffusion model, an intuitive approach emerges: utilizing prompts on the input images to modulate the expression levels of specific features. We term this approach FilterPrompt.
Specifically, we leverage IP-Adapter[50] to obtain the conditional encoding $C_{s}$, which governs the rendering attributes of the generated result from the reference image $x_{style}$. Simultaneously, we employ ControlNet[52] to obtain the conditional encoding $C_{c}$, responsible for regulating the geometric attributes from the reference image $x_{content}$. $C_{s}$ and $C_{c}$ can be regarded as a collection of features from reference images.
$C_{s}=\left\{feature_{\{m\}(x_{\{style\}})}\right\}_{\{m=1\}}^{\{M\}},C_{c}=\left\{feature_{\{n\}(x_{\{content\}})}\right\}_{\{n=1\}}^{\{N\}}$ (3)
However, due to the inability to guarantee disentanglement between the features of $C_{s}$ and $C_{c}$, directly using them for guided generation in the latent space would inevitably result in content conflicts. Drawing upon the successful performance of prompts demonstrated in the dynamic generation model above, we try to alleviate these conflicts by leveraging FilterPrompt. Moreover, features often exhibit distinct differences in frequency or distribution in pixel space. This fact empowers us to tailor the design of FilterPrompt to meet specific task prerequisites. For example, integrating a color-removing prompt $F$ will nullify color features, leading to the subsequent alterations in coding:
$C_{s}^{\prime}=\left\{feature_{\{m\}(F(x_{\{style\}}))}\right\}_{\{m=1\}}^{(M-k)}\quad,C_{c}^{\prime}=\left\{feature_{\{n\}(F(x_{(content)}))}\right\}_{\{n=1\}}^{\{N-k\}}$ (4)
At the same time, the pixel range affected by $F$ is intuitive and convenient, so we can estimate the impact of it on the migration of Gaussian distribution in Equation 2 and change the iterative process into:
@@
In conclusion, we advocate for the adoption of our innovative approach, FilterPrompt, which directly manipulates the frequency or distribution characteristics of specific image attributes, thereby influencing the subsequent expression levels of the representation. This approach helps the diffusion model demonstrate more powerful control capabilities when dealing with complex generation tasks.
4.2 Architecture Description
Framework for Experiments: In order to verify and discuss the impact of our approach on the control effect of diffusion models, our experiment are built based on combined filtering operations, ControlNet [52] and IP-Adapter [50], as shown in Figure 4. This means we combine ControlNet with IP-Adapter as the baseline. Specifically, inspired by the concept of decoupled cross-attention in IP-Adapter, we apply a similar methodology. Subsequently, we map the low-level features in the style image to a global embedding $C_{s}$ and concatenate it with the default text prompt embedding $C_{t}$ of SDM. This process can be described as $X_{t}=C_{t}\bigoplus C_{s}$. These two parts in hidden state $x_{t}$ are processed separately at each denoising generation. A portion of $x_{t}$ is delegated to ControlNet, which projects the latent distribution into a fused distribution controlled by high-level features $C_{c}$. The global embedding of another part in $x_{t}$ utilizes the IP-Adapter for decoding, unfolding, and guiding the generation of low-level features. We use $x_{t-1}$ to represent hidden state predicted at the next moment in Equation 2. The intermediate hidden states obtained from both processes are weighted and summed according to Equation 6, achieving the effect of unifying representations related to Content and Style into the latent space of SDM as shown in Equation 6.
$X_{t-1}=\alpha\cdot ControlNet_{hiddenStates}+\beta\cdot\lambda\cdot IP_{hiddenStates}$ (6)
@@ 4.2 Architecture Description
In experiments 4-2-1 and 4-2-2, $\alpha=1$, $\beta=1$, $\lambda=7.5$, and the random seed of the diffusion model is 42. In experiments 4-2-3, 4-3, and quantitative analysis, $\alpha=1.2$, $\beta=0.6$, $\lambda=7.5$, and the random seed of the diffusion model is 52.
Figure 4: Framework for experiments. The experiment uses ControlNet and IP-Adapter as the baseline and adds combined filtering operations as the expansion. We mapped low-level features in style images to global embeddings as $Cs$, concatenating them with SDM default text prompt embeddings $Ct$. The denoising generation processes these parts separately. A segment is managed by ControlNet, projecting latent distributions into a fused distribution controlled by high-level features that is $Cc$. The other part uses IP-Adapter for decoding and guiding low-level feature generation. Intermediate hidden state $x_{t-1}$ from both processes are weighted and summed every sampling time.
4.3 Effects of Various FilterPrompts on Baseline
For experiments, we define structure to represent the geometric features in the content image and appearance to denote the rendering features derived from the style image’s color distribution and pixel-level texture. Next, we assess and compare the performance variations of the baseline on structural features from content images and appearance features from style images before and after applying FilterPrompt (see Figure 5).
@@ 4.3 Effects of Various FilterPrompts on Baseline
Figure 5: Comparison of the effect of existing approaches and FilterPrompts on structural control. The generated results demonstrate that the combined filters in FilterPrompt(4) have the best preservation effect of structural information. Specifically, these filters enable high-fidelity reproduction of critical vehicular details such as the exhaust window and headlights, which existing preprocessing methodologies fail to replicate.
Firstly, we perform our approach on the ControlNet path, which itself is used for control structure information in the baseline. The comparison of results among existing image preprocessing approaches and FilterPrompt indicates that our approach retains more details in the generated results and brings the colors closer to the ideal appearance. Meanwhile, comparisons between text control and FilterPrompt indicate that the impact of our approach is more pronounced, as shown in Figure 5. In the various FilterPrompt configurations shown in the figure, the combinations (using the ITV-R 601-2 luma transform method for decolorization, inversion, and contrast enhancement of the grayscale image) in Figure 5-FilterPrompt(4) demonstrates the best preservation effect of structural information. The purpose of these filters is to change the brightness distribution of the picture to make the outline structure clearer. Therefore, FilterPrompt exhibits strong intuitiveness and interpretability.
Figure 6: Effect of different FilterPrompts on appearance image. Among the results, the Sharpen filter enhances the expression of fine strokes, while the Gaussian filter blurs detailed stroke information. This illustrates that FilterPrompt can have a discernible impact on appearance information, aligning well with our expectations.
@@
Figure 7: Impact of different kernel sizes in FilterPrompt on the generated results. In this example, we utilize FilterPrompt(4) mentioned before on the structure image. Simultaneously, a Gaussian filter is applied to the appearance image. The outcomes highlight the effectiveness of FilterPrompt(4) on the structure image in preserving the geometric attribute information of the bronze. Additionally, increasing the Gaussian kernel size helps reduce the representation of redundant pattern information in the appearance image, thus addressing content conflicts in the generated results.
@@
Additionally, we also conduct a comparative analysis of the IP-adapter path, which controls appearance information, both before and after the implementation of FilterPrompt. According to the results in Figure 6, we observed that after applying FilterPrompt designed for noise processing to the current appearance image, it significantly affects the detailed stroke information of the generated image. At the same time, the generated effects are consistent with the characteristics of the applied filters themselves. For example, the Sharpen filter enhances the expression of fine strokes, while the Gaussian filter blurs detailed stroke information. This experimental result verifies the capability of FilterPrompt to produce the expected control effects on appearance.
@@
Taking a step further, we explore the control effect of FilterPrompt on both paths in the baseline. In the task that requires high protection of geometric features (Figure 7), the aim is to transfer a bronze sketch to a photo with a specified appearance image without altering geometric features from the structure image. We use the same prompts in Figure 5-FilterPrompt(4) for ControlNet path, to enhance control over structure features. However, the initial results reveal redundant pattern semantic expressions on the bronze surface. So, we apply Gaussian filter for the IP-Adapter path to help suppress high-frequency information. It can be seen from the histogram in Figure 7 that Gaussian filter hardly changes the color distribution but will destroy the continuous distribution of high-frequency information. The generation results show that when the Gaussian kernel increases to 4, the redundant pattern features in the generation results significantly fade.
4.4 Explanation
Figure 8: Sampling inference process. After applying our approach FilterPrompt, the underlying texture in the sampled images changes from a distribution resembling arc patterns to a point-like distribution. Additionally, as shown in the enlarged illustration on the right, it is evident that the use of filters consistently disrupts the expression of redundant pattern features.
We previously discussed that FilterPrompt, when viewed from the perspective of dynamic generation, serves as a guiding force in steering the Gaussian distribution of current data toward the target distribution during migration and diffusion. This controlled influence on distribution migration is elucidated in this subsection by examining the effects of FilterPrompt at various stages during the sampled inference process.
As shown in Figure 8, we perform a detailed analysis of the sampling inference stage for the task of converting a bronze sketch to a photo in the preceding section. Notably, the utilization of Gaussian FilterPrompt interrupts the expression of redundant texture features. The sampled results vividly illustrate that detailed representations of arc patterns initially present in the early stages are weakened by Gaussian FilterPrompt, manifesting as point-like distributions. This observation suggests that Gaussian FilterPrompt disrupts the continuous semantic expression of patterns. Concurrently, subsequent texture generation doesn’t emerge the negative impact of full-graph blurring, so it illustrates the property that FilterPrompt is only effective for specific feature distributions.
The third row of the evolution sequence in Figure 8 illustrates the impact of applying Gaussian FilterPrompt only in the first 12 steps, where we observe the absence of redundant pattern features in the final generation result. A set of comparisons in lines 4 and 5 showcase the effectiveness of Gaussian FilterPrompt. In the first 6 steps, only Controlnrt is utilized to regulate the structural layout. Introducing IP-adapter in the 7th step guides appearance features. The final results reveal that, even if Gaussian FilterPrompt does not initially suppress redundant features, it remains effective in later stages during the generation of detailed textures. The above findings highlight that the impact of filtering on the diffusion process is intuitive, controllable, and predictable.
5 Results
In this part, we qualitatively and quantitatively analyze the control effect impact of our approach by using FilterPrompt(4) in Figure 5. The generated results demonstrate our approach can be conveniently generalized to various data domains and controllable generation tasks.
Figure 9: Comparison with other works.
@@ 5.1 Quantitative analysis
The essence of the appearance transfer task lies in effectively blending the high-level features found in structural images with the low-level features present in appearance images. High-level features in structural images encompass their geometric properties and semantic characteristics, whereas low-level features in appearance images encompass aspects such as color distribution and texture features. Our quantitative analysis focuses on six distinct types of appearance transfer tasks: cat to cat, cat to dog, cat to wild, bird to bird, airplane to bird, and car to car. It’s essential to note that all structure-appearance image pairs utilized in these experiments share the same style: Photographs.
By referring to some works on image transfer tasks [40, 27, 13, 1], we selected six metrics that align with the objectives of our quantitative analysis experiment. To assess the retention of geometric attributes and semantic features from structure images in generated images, we employ three key indicators: Structure Preservation (SP), Chamfer Distance (CD), and Fréchet Inception Distance (FID). Concurrently, to evaluate the fidelity of low-level features between appearance images and generated images, we utilize three distinct metrics: Gray-Level Co-occurrence Matrix (GLCM), Peak Signal-to-Noise Ratio (PSNR), and Color Histogram Correlation (CHC).
@@
Test Datasets: AFHQ [10], CUB-200-2011 [44], FGVC-Airaft [32], Stanford-Cars [24]. Among them, the three domain data of cat, dog, and wild are all from the AFHQ test set. We followed the setting of AFHQ’s test set, with 500 images for each category, and randomly selected 500 images from three other datasets as appearance images. For every types of appearance transfer tasks there are 2000 pairs, Therefore, the data in Table 5 is based on the evaluation results of 12000 Structure-Appearance image pairs.
@@ 5.2 Qualitative analysis
The qualitative analysis experiments include a total of eight domains (cat, dog, wild, bird, airplane, car, bronze, architecture). In addition to the datasets used in the quantitative analysis experiments, additional datasets are: WikiChurches [3] and Bronze Dings [54].
Among them, we mainly designed two types of tasks. One is the style transfer as shown in Figure 10 including three domains (bronze, airplane, architecture). In this task, the style image and the content image do not have semantic correspondence, and their relationships belong to different domains. So the focus in this task is to obtain the low-level texture features from the style image without semantic correspondence and then render it to the content image.
Figure 10: Style transfer tasks. Our focus is to obtain the low-level texture features from the style image without semantic correspondence and then render it to the content image.
The another is the appearance transfer task as shown in Figure 11 that includes six domains (cat, dog, wild, bird, airplane, car). The image pairs used in this task belong to the same data distribution, and focus is to transfer the underlying color distribution and stroke texture in the appearance image to the structure image without destroying its geometric properties.
Figure 11: Appearance transformation tasks. A total of six domains (cat, dog, wild, bird, airplane, car), we showcase the effects achieved by the baseline architecture with filtering combined operation in local, object-centric, and full-image-level appearance transfer tasks.
@@ 6 Limitation and Conclusion
While the field of controllable generation tasks is flourishing, feature decomposition remains a formidable challenge. Therefore, harnessing distinctions in distribution and frequency domains among features in pixel space, we introduce FilterPrompt, an approach designed to amplify the efficacy of model control. This approach boasts universal applicability across any diffusion model, enabling users to adjust the representation of specific image features to task specifications, thereby facilitating more precise and controllable generation outcomes. Through our experiments, we demonstrate that utilizing combinations such as the ITV-R 601-2 luma transform for decolorization, inversion, and contrast enhancement of grayscale images can heighten the diffusion model’s control over structural features, while Gaussian FilterPrompt can mitigate the expression of redundant texture features.
Beyond the baseline framework utilized in this study, the integration of other advanced diffusion models may potentially lead to even superior results if they are implemented to replace certain components in the baseline framework. However, exploring this possibility is beyond the scope of our current discussion. Notably, the design of FilterPrompt requires manual intervention in accordance with the task requirements and data characteristics. This process entails making preliminary judgments regarding the focus of content information and incurring certain trial-and-error costs.
In summary, we contend that the integration of FilterPrompt with diffusion models can achieve remarkably impressive control effects compared to alternative approaches within the feature space. Moreover, it is exceedingly convenient, making it suitable for widespread adoption in other diffusion models.

### v3

@@ 4.1 FilterPrompt
We leverage IP-Adapter[56] to obtain conditional encoding $C_{s}$, which controls rendering attributes in the generated image from the appearance reference $x_{app}$. Concurrently, we employ ControlNet[58] to obtain the conditional encoding $C_{c}$, responsible for regulating the geometric attributes from the structure reference image $x_{struct}$. These encodings, $C_{s}$ and $C_{c}$, represent features from their respective reference images:
$C_{s}=\left\{feat_{\{m\}(x_{\{app\}})}\right\}_{\{m=1\}}^{\{M\}}$ (3)
$C_{c}=\left\{feat_{\{n\}(x_{\{struct\}})}\right\}_{\{n=1\}}^{\{N\}}$
However, since $C_{s}$ and $C_{c}$ features are not fully disentangled, using them directly in latent space may cause structural conflicts. Drawing upon the successful performance of prompts demonstrated in the dynamic generation model above, we try to alleviate these conflicts by leveraging FilterPrompt, which adjusts feature frequencies or distributions. For instance, a color-removing prompt $F$ removes color features $feat$, modifying the encodings as follows:
$C_{s}^{\prime}=\left\{feat_{\{m\}(F(x_{\{app\}}))}\right\}_{\{m=1\}}^{(M-k)}$ (4)
$C_{c}^{\prime}=\left\{feat_{\{n\}(F(x_{(struct)}))}\right\}_{\{n=1\}}^{\{N-k\}}$
The pixel range affected by $F$ is intuitive and convenient, so we can estimate the impact of it on the migration of Gaussian distribution in Equation 2 and change the iterative process into:
@@
In conclusion, we advocate for the adoption of our innovative approach, FilterPrompt, which directly manipulates the frequency or distribution characteristics of specific image attributes, thereby influencing the subsequent expression levels of the representation.
4.2 Architecture Details
Our framework are built based on combined filtering operations, ControlNet [58] and IP-Adapter [56], as shown in Figure 5. Subsequently, we map the low-level features in the appearance image to a global embedding $C_{s}$ and concatenate it with the default text prompt embedding $C_{t}$ of SDM. This process can be described as $X_{t}=C_{t}\bigoplus C_{s}$. These two parts in hidden state $x_{t}$ are processed separately at each denoising generation. A portion of $x_{t}$ is delegated to ControlNet, which projects the latent distribution into a fused distribution controlled by high-level features $C_{c}$. The global embedding of another part in $x_{t}$ utilizes the IP-Adapter for decoding, unfolding, and guiding the generation of low-level features. We use $x_{t-1}$ to represent the hidden state predicted at the next moment in Equation 2. The intermediate hidden states obtained from both processes are weighted and summed according to Equation 6, achieving the effect of unifying representations related to Structure and Appearance into the latent space of SDM as shown in Equation 6.
$X_{t-1}=\alpha\cdot ControlNet+\beta\cdot\lambda\cdot IP-Adapter$ (6)
@@ 4.2 Architecture Details
Figure 6: Figure-left: The effect of FilterPrompt applied to the appearance image. Among the results, the Sharpen filter enhances the expression of fine strokes, while the Gaussian filter blurs detailed stroke information. This demonstrates that FilterPrompt can significantly influence appearance information, aligning with our expectations. Figure-right: The effect of FilterPrompt applied to the structure image. The generated results show that the $FilterPrompt_{struct}$ in FilterPrompt(4) best preserve structural information. Specifically, these filters allow high-fidelity reproduction of critical vehicular details, such as the exhaust window and headlights, which existing preprocessing methods fail to replicate.
@@
Figure 7: Impact of different kernel sizes in FilterPrompt on the generated results. In this example, we utilize $FilterPrompt_{struct}$ mentioned before on the structure image. Simultaneously, a Gaussian filter is applied to the appearance image. The outcomes highlight the effectiveness of $FilterPrompt_{struct}$ on the structure image in preserving the geometric attribute information of the bronze. Additionally, increasing the Gaussian kernel size helps reduce the representation of redundant pattern information in the appearance image, thus addressing content conflicts in the generated results.
4.3 Effects of Various FilterPrompts
We define structure as the geometric features in the structure image and appearance as the rendering features from the reference image’s color and texture. We then assess baseline performance on these features before and after applying FilterPrompt (see Figure 6).
Firstly, we apply our approach on the ControlNet path for controlling structural details. The comparison of results among existing image preprocessing approaches and FilterPrompt indicates that our approach retains more details in the generated results and brings the colors closer to the target appearance. The combination of operations (including the ITV-R 601-2 luma transform method for decolorization, inversion, and contrast enhancement of the grayscale image) is succinctly referred to as $FilterPrompt_{struct}$. As illustrated in Figure 6-FilterPrompt(4), $FilterPrompt_{struct}$ best preserves structural information by enhancing brightness for clearer outlines, demonstrating FilterPrompt’s clarity and interpretability.
@@ 4.3 Effects of Various FilterPrompts
We then analyze the IP-adapter path, responsible for appearance, before and after applying FilterPrompt. Figure 6 shows that applying noise-processing filters to the appearance image affects the generated stroke details, aligning with the filter’s effects—e.g., Sharpen enhances fine strokes, while Gaussian blurs them. This verifies FilterPrompt’s capacity to control appearance.
@@
Taking a step further, we explore the control effect of FilterPrompt on both paths in the baseline. In appearance task as shown in Figure 7), we aim is to transfer a bronze sketch to a photo with a specified appearance image without altering geometric features from the structure image. For structure control, $FilterPrompt_{struct}$ in the ControlNet path was used, but initial results showed redundant patterns. Applying a Gaussian filter in the IP-Adapter path suppressed high-frequency noise. When the Gaussian kernel increased to 4, redundant features diminished significantly.
5 Experimental Results
@@
Figure 8: Comparison with other works. As shown in the 1st, 2nd, 3rd, 5th, and 8th columns of the figure, the baseline has content conflict problems, the Cross-image attention generates blurry results, and the IP-Adapter cannot accurately transfer the color of the appearance image. Applying $FilterPrompt_{struct}$ on the baseline results can enhance the protection of structural attributes while alleviating content conflicts.
Figure 9: Appearance transfer tasks. We showcase the effects achieved by the baseline architecture with filtering combined operation in appearance transfer tasks.
@@ 5.1 Quantitative analysis
Our quantitative analysis covers six specific appearance transfer: cat to cat, cat to dog, cat to wild, bird to bird, airplane to bird, and car to car. Following previous works [44, 30, 13, 1], we selected six metrics for our experiment. To evaluate the retention of geometric and semantic features from structure images in the generated images, we use three primary indicators: Structure Preservation (SP), Chamfer Distance (CD), and Fréchet Inception Distance (FID). Additionally, to assess the fidelity of low-level features between appearance images and generated images, we employ three specific metrics: Gray-Level Co-occurrence Matrix (GLCM), Peak Signal-to-Noise Ratio (PSNR), and Color Histogram Correlation (CHC).
@@
Test Datasets: AFHQ [10], CUB-200-2011 [49], FGVC-Airaft [35], Stanford-Cars [26]. Among them, the three domain data of cat, dog, and wild are all from the AFHQ test set. We followed the setting of AFHQ’s test set, with 500 images for each category, and randomly selected 500 images from three other datasets as appearance images. For every types of appearance transfer tasks there are 2000 pairs, Therefore, the data in Table 5 is based on the evaluation results of 12000 Structure-Appearance image pairs. We show examples of the comparison results in Figure 8.
@@ 5.2 Qualitative analysis
The qualitative analysis experiments include a total of five domains (cat, dog, wild, bird, bronze). In addition to the datasets used in the quantitative analysis experiments as shown in Figure 9, additional data is: Bronze Dings [60]. In this task, the appearance reference and the structure image do not have semantic correspondence, and their relationships belong to different domains. So the focus in this task is to obtain the low-level texture features from the appearance image without semantic correspondence and then render it to the structure image.
5.3 User Study
We conducted a user study with 18 questions evaluating generated results in structure preservation, rendering feature transfer effectiveness, and overall quality. Each participant was paid 0.5\$. To address the common issue of low-quality feedback caused by participants’ lack of understanding of task, we had eight facilitators provide detailed background information to participants on campus. Participants were then asked to select the most fitting options from anonymous choices based on their preferences. Finally, we received 215 valid survey submissions, with ours garnering a support rate of 51. 89% (cross-image 24. 00%, IP-Adapter 14. 98%, ControlNet + IP-Adapter 9. 13%).
@@ 6 Limitation and Conclusion
We constructed an experimental framework based on IP-Adapter to explore image generation techniques. However, we found that the identity consistency of the generated results is not always satisfactory when the input image prompt contains rich semantic information. This limitation is partly because IP-Adapter may not fully balance these properties when processing different image attributes, thus affecting the consistency of the final output, which may be attempted by using mask technology to optimize identity consistency and image quality. Beyond the baseline framework utilized in this study, the integration of other advanced diffusion models may potentially lead to even superior results if they are implemented to replace certain components in the baseline framework. We will explore this further in future work.
In conclusion, we propose a pixel-space processing, FilterPrompt, to guide image appearance transfer, by focusing on the input sensitivity and dynamic evolution of diffusion models. We find that the models adapt to input feature distributions, enabling targeted operations for precise control in final generated images. Experimental results show that our approach helps refine structural control and reduce redundant textures in transfer tasks. Although FilterPrompt requires manual setup, it provides a simple yet efficient way to enhance customization and control in diffusion models.